Five years ago, Brad was two years older than 5/6 of his age today. What is his age today
Answer by oberobic(2304)

Brad's current age is: x
5 years ago, Brad's age was: x-5
At that time, his age was 2 more than 5/6*x.
What is x?
.
x-5 = 5/6*x + 2
.
Multiply through by 6 to eliminate the fraction
6x - 30 = 5x + 12
.
Subtract 5x from both sides
x - 30 = 12
.
Add 30 to both sides
.
x = 42
.
Five years ago, Brade's age was: x-5 = 37
5/6 of his current age is: 5/6x = 35
which is 2 less than 37.
Done